OpenStudy (ny,ny):

(sqrt8 + sqrt12)^2 + (sqrt8 - sqrt12)^2 = c^2

OpenStudy (ny,ny):

c is hypotenuse. its a triangle question.

OpenStudy (dumbcow):

FOIL

OpenStudy (anonymous):

Damn forget I said anything

OpenStudy (ny,ny):

really.

OpenStudy (ny,ny):

... i know i have to foil it and then simplify like write each set of parenthesis twice so after writing all that out can i do (sqrt64+sqrt96+sqrt96+sqrt144)+(sqrt64-sqrt96-sqrt96+sqrt144) or is that wrong.

OpenStudy (anonymous):

\[\left(2 \sqrt{2}+2 \sqrt{3}\right)^2+\left(2 \sqrt{2}-2 \sqrt{3}\right)^2=c^2 \]\[c^2=40 \]\[c=2 \sqrt{10},c=-2 \sqrt{10} \]

OpenStudy (ny,ny):

thank you.